### v1

@@
Abstract
With the release of ChatGPT, OpenAI has reinforced its leadership in large language models. Simultaneously, OpenAI offers developers advanced technologies and resources (such as API services), driving innovation and development in various applications. This initiative promotes technological advancement and inspires developers to further explore and address the challenges of applying these technologies.
@@ Abstract
In contrast to traditional software development practices, OpenAI’s development introduces new challenges for AI developers in design, implementation, and deployment. These challenges span different areas (such as prompts, plugins, and APIs), requiring developers to navigate unique methodologies and considerations specific to large language model development.
@@
However, to the best of our knowledge, these challenges have not been thoroughly investigated in previous empirical studies. To fill this gap, we present the first comprehensive study on understanding the challenges faced by OpenAI developers. Specifically, we crawl and analyze 29,057 relevant questions from a popular OpenAI developer forum. We first examine their popularity and difficulty.
After manually analyzing 2,364 sampled questions, we construct a taxonomy of challenges faced by OpenAI developers. Based on this taxonomy, we summarize a set of findings and actionable implications for OpenAI-related stakeholders, including developers and the OpenAI organization.
Keywords:
@@
Mining Software Repository, Empirical Study, OpenAI Developer, Development Challenges, Prompt Engineering
journal: Empirical Software Engineering
1 Introduction
The emerging large language models (LLMs) are increasingly attracting attention and becoming one of the hot research topics in the field of computer science. Until now, LLMs have demonstrated promising performance in different fields, such as natural language processing and software engineering (Zhao et al, 2023b; Wang et al, 2024; Fan et al, 2023).
OpenAI, an artificial intelligence research organization, has been at the forefront of developing large language models (LLMs), which are AI systems designed to understand and generate human-like text. These models, such as GPT-3 and its successor GPT-4, have garnered significant attention for their ability to perform a wide range of language-based tasks with impressive performance.
With the rise in popularity of ChatGPT and the rapid development of related OpenAI technologies, an increasing number of developers are utilizing these tools to assist in their development processes.
However, during OpenAI development, developers often encounter various challenges. For instance, correctly configuring and invoking OpenAI’s API can be difficult, including setting parameters, managing rate limits, and handling errors. For those unfamiliar with AI and LLMs, developing plugins and applications based on OpenAI’s technology can be daunting, involving integration, performance optimization, and ensuring security. Ensuring data privacy and security while handling user data is crucial. Developers must comply with relevant regulations and implement necessary security measures.
@@ 1 Introduction
Compared to traditional software development, OpenAI developers face the following unique challenges: (1) Automating task processing: OpenAI’s large language models can automatically handle many tasks, such as text generation, image recognition, and speech recognition. In contrast, traditional software engineering often requires manual coding of various algorithms and logic to accomplish these tasks. (2) Dealing with uncertainty: LLMs can produce variable and sometimes unpredictable outputs, unlike traditional software which typically has deterministic outcomes. Developers must account for this uncertainty and implement mechanisms to manage it. (3) Handling large-scale datasets: OpenAI development often involves working with large-scale datasets to train or fine-tune models. This requires specialized knowledge in data preprocessing, management, and the efficient use of computational resources. (4) Data privacy and security: Using LLMs requires a significant amount of data for training and fine-tuning, raising concerns about the privacy and security of user data. Developers need to design effective methods to ensure data security and privacy. (5) Performance optimization: Optimizing the performance of large language models (LLMs), such as output accuracy, differs from optimizing traditional software. (6) Interpreting model outputs: Understanding and interpreting the outputs of LLMs can be complex. Developers need to ensure that the model’s outputs are reliable and contextually appropriate. These challenges require a different set of skills and considerations, highlighting the distinct nature of OpenAI development compared to traditional software development. However, to the best of our knowledge, these challenges faced by OpenAI developers have not been thoroughly investigated in previous empirical studies.
To fill this gap, we present the first comprehensive study on understanding the challenges for OpenAI developers. Our empirical study aims to help developers to avoid common pitfalls and improve development efficiency. To this end, we crawl and analyze 29,057 questions from the OpenAI developer forum [^1] https://community.openai.com.
@@
The OpenAI Developer Forum is a collaborative space for developers to seek assistance and share insights on using OpenAI technologies. It supports developers of all skill levels, offering discussions on API integration, plugin development, and best practices. The forum also provides updates on the latest advancements and tools from OpenAI.
@@
RQ1: What is the popularity trend of OpenAI development among developers?
Result. After analyzing posts related to OpenAI development, we examine the number of new posts and new users added over each period. Our findings indicate that OpenAI development is attracting increasing attention from developers, particularly with the introduction of pivotal products like ChatGPT. This growing trend underscores the rising popularity of OpenAI technologies, highlighting the timeliness and importance of our empirical study.
RQ2: How difficult is OpenAI development for developers?
Result. To illustrate the difficulty level of the challenges faced by OpenAI developers, we analyze the number of replies to OpenAI development-related questions. The results show that 53% of these questions have fewer than three replies, suggesting that they are typically challenging to address. This finding prompts us to further investigate the underlying challenges of these development questions.
RQ3: What specific challenges for OpenAI developers?
Result. We perform a manual analysis on 2,364 sampled questions and construct a taxonomy of challenges consisting of 27 categories. For example, Prompt Design, Integration with Custom Applications, and Token Limitation. In addition, based on this taxonomy, we summarize findings and actionable implications for stakeholders (such as developers and the OpenAI organization).
The main contributions of our empirical study can be summarized as follows:
  • We conduct the first empirical study to investigate the challenges faced by OpenAI developers by mining the OpenAI developer forum.
@@
Paper Organization. The rest of this paper is organized as follows. Section 2 introduces the background of OpenAI’s development. Section 3 describes our methods for collecting and analyzing forum posts. Section 4 analyzes the popularity of the forum. Section 5 analyzes the difficulty levels of the posts in the forum. Section 6 details our constructed taxonomy and discusses related insights. Section 7 describes validity threats. Section 8 discusses work related to development challenges. Section 9 summarizes this research.
2 Background
@@
OpenAI has emerged as a leading organization in artificial intelligence research, offering a range of APIs that enable developers to integrate advanced AI capabilities into their applications. The OpenAI API provides access to various models that can perform a wide array of tasks, from natural language understanding and generation to more complex functions like code completion and image recognition. This section provides an overview of the API usage and the development of plugins within the OpenAI ecosystem.
@@
OpenAI Usage. The OpenAI API allows developers to leverage the power of state-of-the-art AI models through a simple interface. Key features of the API include: (1) Ease of Integration: Developers can integrate the API into their applications with minimal setup, using familiar HTTP requests. (2) Versatile Capabilities: The API supports various tasks including text generation, translation, summarization, and more. (3) Scalability: The API is designed to handle a wide range of requests, from small-scale personal projects to large-scale enterprise applications. The API offers several endpoints corresponding to different models and capabilities. For instance, the davinci model is known for its high performance in generating coherent and contextually relevant text, making it suitable for applications requiring sophisticated language understanding.
@@
Plugin Development. Plugins are extensions that enhance the functionality of software applications by embedding AI capabilities directly within them. OpenAI provides robust support for plugin development, allowing developers to create custom plugins that interact with OpenAI models. Key aspects of plugin development include: (1) API Integration: Plugins typically make use of the OpenAI API to fetch responses from AI models based on user input or other triggers. (2) Custom Functionality: Developers can tailor the behavior of plugins to meet specific needs, such as automating customer support responses, generating content, or providing real-time data analysis. (3) Interoperability: OpenAI plugins can be integrated with various platforms and services, enhancing the AI capabilities of a wide range of applications from web and mobile apps to enterprise systems.
@@
OpenAI has introduced a customized version of ChatGPT called GPTs [^3] https://openai.com/index/introducing-gpts. Similar to plugins, GPTs serve personalized user needs. Developers can create different GPTs for specific purposes, which can be shared with others. GPTs offer a new way to create a tailored version of ChatGPT, assisting in daily life and specific tasks. For instance, GPTs can assist in analyzing math problems, learning soccer rules, or even helping in designing icons. Moreover, creating GPTs is a simple process that does not require coding, allowing anyone to easily build their GPTs by engaging in conversation with ChatGPT, providing instructions and additional knowledge, and selecting actions like web searching, data analysis, or image creation.
@@
While the integration of OpenAI’s APIs and the development of plugins and GPTs offer significant advantages, they also present several challenges, such as: (1) Cost Management: Managing the computational and financial costs associated with training, fine-tuning, and deploying large AI models, which can be significantly higher than those for traditional software systems. (2) Ethical Considerations: Ensuring that AI usage adheres to ethical guidelines, particularly in areas like content generation and decision-making. (3) Output authenticity: Verifying the accuracy and reliability of the output generated by artificial intelligence models to ensure that plugins and GPTs meet development goals and user expectations.
3 Methodology
@@
To understand the challenges faced by OpenAI developers, we collect and analyze relevant question posts from the OpenAI developer forum, following previous studies (Ahmed and Bagherzadeh, 2018; Alshangiti et al, 2019; Bagherzadeh and Khatchadourian, 2019; Lou et al, 2020; Rosen and Shihab, 2016; Yang et al, 2016; Zhang et al, 2019). Fig. 1 provides a methodology overview of our empirical study.
Figure 1: Methodology overview of our empirical study.
@@
Step 1: Crawling question posts. To analyze the popularity trends and difficulty levels of the questions, we crawl posts from the OpenAI developer forum. To analyze challenges faced by OpenAI developers, we resort to categories, which make it easier to find relevant posts and ensure that questions reach the appropriate experts. Our study focuses on four specific categories (i.e., API, Prompting, GPT builders, and ChatGPT). Specifically, posts in the “API” category concern questions, feedback, and best practices around building with OpenAI’s API. posts in the “Prompting” category learn more about prompting by sharing best practices, and favorite prompts. GPTs are a new way for anyone to create a tailored version of ChatGPT to be more helpful in their daily life, therefore, the “GPT builders” category is designed for people who are building GPTs. Finally, the “ChatGPT” concerns questions or discussions about ChatGPT. since posts in the remaining categories (i.e., Community, Announcements, Documentation, and Forum feedback) rarely discuss the challenges faced by OpenAI developers, we ignore these posts in our empirical study. Specifically, posts in the “Announcements” category typically contain official product updates, those in the “Documentation” category offer tutorials and instructional content, posts in the “Community” category involve project and technology sharing, and posts in the “Forum feedback” category focus on improving the developer forum.
@@
Fig. 2 shows a screenshot of a post [^4] https://community.openai.com/t/547685 from the OpenAI developer forum. Typically, a post consists of several key elements: title, category, tags, question description, and code snippet, with a few posts having accepted answers. These posts also include additional relevant information, such as creation time, last reply time, reply count, view count, and the number of participating users. In this step, we gather the following metadata: the post’s title, creation date, and the number of replies. We also collect user metadata including usernames and registration times. In summary, by June 2024, we had crawled data from 29,057 posts and 722,389 users.
Figure 2: A question post from the OpenAI developer forum.
Step 2: Determining popularity trend.
@@
To analyze the popularity trend of OpenAI development among developers, we first perform time series analysis. Specifically, we count the number of new posts and new users over different periods (i.e., every three months) from the forum’s inception in February 2021 until our data collection deadline in June 2024. Detailed result analysis can be found in Section 4.
Step 3: Determining difficulty level. In our primary investigation, we find that a small proportion of posts mark one of the replies as “Solution” (i.e., indicating the acceptance by the questioner). The possible reason is that users in this community do not strictly adhere to this practice, unlike those in Stack Overflow (Yazdaninia et al, 2021). Therefore, the methodology of determining difficulty levels through the proportion of accepted answers, as employed in prior studies (Wen et al, 2021; Wang et al, 2023a; Chen et al, 2020), is not applicable in our study. In our study, we mainly assess the difficulty level of questions solely based on the number of replies received. Our study assumes that a lower number of replies indicates a higher level of difficulty in addressing the question. Detailed result analysis can be found in Section 5.
@@
Step 4: Constructing the taxonomy of challenges. Due to the high cost of manually analyzing all the crawled posts, we randomly select a statistically significant sample ensuring a 99% confidence level with a $\pm$2.5confidence interval from 29,057 questions (Aghajani et al, 2019) and remove posts that have already been deleted by OpenAI. As a result, 2,364 questions constitute the taxonomy dataset of our study. We construct the taxonomy as follows.
First, we randomly select 70% of the 2,364 posts to construct the initial taxonomy of challenges by following the previous study (Wen et al, 2021). This process involves two annotators working collaboratively. We use the open coding procedure (Seaman, 1999) to analyze the sampled posts, summarize the preliminary taxonomy, and construct a multi-level hierarchical taxonomy of challenges faced by OpenAI developers. Each annotator has two years of experience in OpenAI development and three years of experience in developing large language models (LLMs). They carefully read the questions of the sampled posts to understand the specific challenges faced by OpenAI developers and the corresponding replies. The open coding procedure is conducted as follows: the two annotators read all the sampled posts, with each post being read at least twice. During the analysis, they examine all the data contained in each post, including titles, problem descriptions, code snippets, and replies, summarizing the specific challenges discussed. Then they summarize the specific challenges discussed in the posts. Specifically, some posts describe issues encountered during the use of ChatGPT, for example, “I can’t login to ChatGPT - what should I do?”  [^5] https://community.openai.com/t/188817, which are not closely related to developers.
@@
In addition, there are also some questions related to account problems, tips for users, etc. These posts are labeled as “Unrelated”. Since the OpenAI developer forum is a platform for all OpenAI users to communicate, some posts are official announcements, document introductions, project discussions, and project sharing. For example, “I have an idea about creating a mix of emoji”  [^6] https://community.openai.com/t/22180, these posts are temporarily labeled as “Unrelated”. If a post is labeled as “Unrelated”, it means that these posts are not considered when constructing the taxonomy because they are not closely related to OpenAI developers. For the remaining posts, the annotators use simple phrases to summarize the developer’s challenges during the first reading. Specifically, some problem descriptions are about developers consulting on parameter settings when calling OpenAI’s APIs, for example, “GPT-4-Vision-Preview fidelity/detail-parameter”  [^7] https://community.openai.com/t/477563, the annotators find that this problem is asked because developers do not know how to set API parameters. Therefore, for these issues, the annotators can temporarily describe them as “API Usage Issue, Parameter Setting”.
@@
Finally, we refine the taxonomy of challenges and conduct a reliability analysis. Specifically, the two annotators independently label the remaining 30% of the posts based on the generated preliminary taxonomy. They label each post as “Unrelated” or assign it to the corresponding category. If some posts can not be categorized, they are temporarily labeled as “Pending”. We use Cohen’s Kappa (k(Cohen, 1960) to measure the consistency between the two annotators during the independent labeling process. The computed $k$ value is 0.812, indicating that the two annotators are almost completely agree (Landis and Koch, 1977). This demonstrates the reliability of our open coding procedure. In addition, for posts labeled as “Pending”, we introduce the third arbitrator who discusses and analyzes the posts with the two annotators to determine which category these posts ultimately belong to. If they can not be categorized into the preliminary taxonomy, new categories are added. In this phase, we incorporate nine additional leaf categories into the preliminary taxonomy framework. Finally, our constructed taxonomy covers all posts related to OpenAI development, and all participants agree on the classification. Detailed result analysis can be found in Section 6.
4 RQ1: Popularity Trend Analysis

### v3

@@
Abstract.
@@
In recent years, large language models (LLMs) have seen rapid advancements, significantly impacting various fields such as computer vision, natural language processing, and software engineering. These LLMs, exemplified by OpenAI’s ChatGPT, have revolutionized the way we approach language understanding and generation tasks. However, in contrast to traditional software development practices, LLM development introduces new challenges for AI developers in design, implementation, and deployment. These challenges span different areas (such as prompts, APIs, and plugins), requiring developers to navigate unique methodologies and considerations specific to LLM application development.
@@
Despite the profound influence of LLMs, to the best of our knowledge, these challenges have not been thoroughly investigated in previous empirical studies. To fill this gap, we present the first comprehensive study on understanding the challenges faced by LLM developers. Specifically, we crawl and analyze 29,057 relevant questions from a popular OpenAI developer forum. We first examine their popularity and difficulty. After manually analyzing 2,364 sampled questions, we construct a taxonomy of challenges faced by LLM developers. Based on this taxonomy, we summarize a set of findings and actionable implications for LLM-related stakeholders, including developers and providers (especially the OpenAI organization).
@@
Mining Software Repository, Empirical Study, LLM Developer, Development Challenges, Prompt Engineering
copyright: acmlicensedjournal: TOSEMjournalyear: 2023ccs: Software and its engineering Software creation and managementccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Emerging large language models (LLMs) are increasingly attracting attention and becoming hot research topics in computer science. Until now, LLMs have demonstrated promising performance in different fields, such as computer vision, natural language processing, and software engineering (Zhao et al., 2023; Wang et al., 2024; Fan et al., 2023). With the rise in popularity of ChatGPT developed by the OpenAI organization and the rapid development of related LLM technologies (Aydın and Karaarslan, 2023; Nazir and Wang, 2023), an increasing number of developers are utilizing these techniques to assist in their development processes. However, during LLM development, developers often encounter various challenges. For example, correctly configuring and invoking LLM’s API can be difficult, including setting parameters, managing rate limits, and handling errors. Developing plugins and applications can be daunting for those unfamiliar with AI and LLMs, involving integration, performance optimization, and ensuring security. Ensuring data privacy and security while handling user data is crucial. Developers must comply with relevant regulations and implement necessary security measures. Therefore, compared to traditional software development, LLM developers face the following unique challenges: (1) LLMs can automatically handle many tasks, such as text generation, image recognition, and speech recognition. In contrast, traditional software engineering often requires manual coding of various algorithms and logic to accomplish these tasks. (2) LLMs can produce variable and sometimes unpredictable outputs, unlike traditional software which typically has deterministic outcomes. Developers must account for this uncertainty and implement mechanisms to manage it. (3) LLM development often involves working with large-scale datasets to train or fine-tune models. This requires specialized knowledge in data preprocessing, management, and the efficient use of computational resources. (4) Using LLMs requires a significant amount of data for training and fine-tuning, raising concerns about the privacy and security of user data. Developers need to design effective methods to ensure data security and privacy. (5) Optimizing the performance of LLMs, such as output accuracy, differs from optimizing traditional software. (6) Understanding and interpreting the outputs of LLMs can be complex. Developers need to ensure that the model’s outputs are reliable and contextually appropriate. These challenges require a different set of skills and considerations, highlighting the distinct nature of LLM application development compared to traditional software development. However, to the best of our knowledge, these challenges faced by LLM application developers have not been thoroughly investigated in previous empirical studies.
@@
To fill this gap, we present the first comprehensive study on understanding the challenges for LLM developers. Our empirical study aims to help developers to avoid common pitfalls and improve development efficiency. To this end, we crawl and analyze 29,057 questions from the OpenAI developer forum [^1] https://community.openai.com. The OpenAI developer forum is a collaborative space for LLM developers to seek assistance and share insights on using OpenAI technologies. It supports developers of all skill levels and offers discussions on API integration, plugin development, and best practices. The forum also provides updates on the latest advancements and tools from OpenAI.
@@
RQ1: What is the popularity trend of LLM development among developers?
Result. After analyzing posts related to LLM development, we examine the number of new posts and new users added over each period. Our findings indicate that LLM development is attracting increasing attention from developers, particularly with the introduction of OpenAI pivotal products like ChatGPT. This growing trend underscores the rising popularity of LLM technologies, highlighting the timeliness and importance of our empirical study.
RQ2: How difficult is LLM development for developers?
Result.
To comprehensively illustrate the difficulty level of the challenges faced by LLM developers, we analyze the number of replies to LLM development-related questions, the proportion of posts with an accepted answer, and the average time required to receive the first reply. The results show that 54% of the questions receive fewer than three replies, only 8.98% of posts have an accepted answer, and it takes an average of at least 147 hours (about six days) to receive the first reply. These findings indicate that questions faced by LLM developers are often difficult to resolve.
RQ3: What specific challenges for LLM developers?
Result. We perform a manual analysis on 2,364 sampled questions and construct a taxonomy of challenges consisting of 26 categories. For example, Prompt Design, Integration with Custom Applications, and Token Limitation. In addition, based on this taxonomy, we summarize findings and actionable implications for LLM stakeholders (such as developers and providers). By comparing our findings to previous empirical studies on ML (Machine Learning)/DL (Deep Learning) developers (Alshangiti et al., 2019; Morovati et al., 2024; Zhang et al., 2019), we reveal that LLM developers encounter unique challenges. For example, prompt design plays a critical role, as developers must carefully craft and refine prompts to optimize LLM output, a process that differs greatly from traditional ML/DL’s focus on feature engineering. Additionally, LLMs often exhibit hallucinations, generating incorrect or non-existent information, which is a less frequent issue in ML/DL models. Another significant challenge is the reproducibility of results, where uncertainty in the generation process or LLM updates can lead to inconsistent outputs. Lastly, the high cost of API calls adds a significant financial burden, a challenge not commonly encountered in traditional ML/DL workflows.
Though LLM developer challenges are primarily analyzed by mining the OpenAI developer forum, our study has significant implications beyond the scope of OpenAI’s LLMs. To verify this, we reconstruct the challenge taxonomy by analyzing issues related to more LLMs (such as Llama provided by Meta AI and Gemini provided by Google) on GitHub. Therefore, our study’s findings broadly apply to LLM products developed by other companies. The commonalities in the challenges developers face (such as prompt design, the hallucination problem, and result reproducibility) are not unique to OpenAI. Still, they are prevalent across the entire LLM ecosystem. Therefore, the insights gained from our research can inform best practices, guide the development of support tools, and improve documentation and community resources for developers working with any LLM-based applications.
To the best of our knowledge, the main contributions of our empirical study can be summarized as follows:
  • We conduct the first empirical study to investigate the challenges faced by LLM application developers by mining the OpenAI developer forum.
@@
Paper Organization. The rest of this paper is organized as follows. Section 2 introduces the background of LLM development. Section 3 describes our methods for collecting and analyzing forum posts. Section 4 analyzes the popularity of the forum. Section 5 analyzes the difficulty levels of the posts in the forum. Section 6 details our constructed taxonomy and discusses related findings. In Section 7, we first discuss the generalization of the challenge taxonomy by analyzing issues to more LLMs on GitHub. Then we analyze potential threats and the efforts to alleviate these threats. Section 8 discusses work related to our study. Section 9 summarizes our empirical study.
2. Background
@@
The development of large language models has revolutionized the field of computer vision, natural language processing, and even software engineering (Zhao et al., 2023; Wang et al., 2024; Fan et al., 2023). As a leading organization in LLM research, OpenAI offers a range of APIs that enable developers to integrate advanced LLM capabilities into their applications. The OpenAI API provides access to various LLMs that can perform a wide array of tasks, from natural language understanding and generation to more complex tasks like code generation/completion and image recognition. This section provides an overview of the API usage and the development of plugins within the OpenAI ecosystem.
@@
API Service. The API service of the LLM allows developers to leverage the power of state-of-the-art LLMs through a simple interface. Key features of the API include: (1) Developers can integrate the API into their applications with minimal setup, using familiar HTTP requests. (2) The API supports various tasks including text generation, translation, summarization, and more. (3) The API is designed to handle a wide range of requests, from small-scale personal projects to large-scale enterprise applications. The API offers several endpoints corresponding to different models and capabilities. For instance, the davinci model is known for its high performance in generating coherent and contextually relevant text, making it suitable for applications requiring sophisticated language understanding.
@@
Plugin Development. Plugins are extensions that enhance the functionality of software applications by embedding AI capabilities directly within them. For example, OpenAI provides robust support for plugin development, allowing developers to create custom plugins that interact with models provided by OpenAI. Key aspects of plugin development include: (1) Plugins typically make use of the OpenAI API to fetch responses from AI models based on user input or other triggers. (2) Developers can tailor the behavior of plugins to meet specific needs, such as automating customer support responses, generating content, or providing real-time data analysis. (3) OpenAI plugins can be integrated with various platforms and services, enhancing the AI capabilities of a wide range of applications from web and mobile apps to enterprise systems.
@@
In November 2023, OpenAI introduced a customized version of ChatGPT called GPTs [^3] https://openai.com/index/introducing-gpts and developed GPT Store [^4] https://openai.com/index/introducing-the-gpt-store/ for sharing developed GPTs. Similar to plugins, GPTs serve personalized user needs. Developers can create different GPTs for specific purposes, which can be shared with others. GPTs offer a new way to create a tailored version of ChatGPT, assisting in daily life and specific tasks. For instance, GPTs can assist in analyzing math problems, learning soccer rules, or even helping in designing icons. Moreover, creating GPTs is a simple process that does not require coding, allowing anyone to easily build their GPTs by engaging in conversation with ChatGPT, providing instructions and additional knowledge, and selecting actions like web searching, data analysis, or image creation.
@@
While the integration of LLM’s APIs and the development of plugins and GPTs offer significant advantages, they also present several challenges, such as: (1) Managing the computational and financial costs associated with training, fine-tuning, and deploying LLM models, which can be significantly higher than those for traditional software systems. (2) Ensuring that LLM usage adheres to ethical guidelines, particularly in areas like content generation and decision-making. (3) Verifying the accuracy and reliability of the output generated by LLMs to ensure that plugins and GPTs meet development goals and user expectations.
3. Methodology
@@
To understand the challenges faced by LLM developers, we collect and analyze relevant question posts from the OpenAI developer forum, which is a forum provided by OpenAI for developers working with their LLM products. We mined this forum by following previous studies (Ahmed and Bagherzadeh, 2018; Alshangiti et al., 2019; Bagherzadeh and Khatchadourian, 2019; Lou et al., 2020; Rosen and Shihab, 2016; Yang et al., 2016; Zhang et al., 2019). Fig. 1 provides a methodology overview of our empirical study.
Figure 1. Methodology overview of our empirical study.
@@
Step 1: Crawl relevant data from the forum. To analyze the popularity trends and difficulty levels of the questions, we crawl posts from the OpenAI developer forum. To analyze challenges faced by LLM developers, we select related subforums of this forum (each subforum’s posts are centered around a specific topic, which makes it easier to find relevant posts and ensures that questions reach the appropriate experts). Our study focuses on four subforums (i.e., API, Prompting, GPT builders, and ChatGPT) in this forum. Specifically, posts in the “API” subforum concern questions, feedback, and best practices around building with OpenAI’s product API. posts in the “Prompting” subforum learn more about prompting by sharing best practices, and favorite prompts. GPTs are a new way for anyone to create a tailored version of ChatGPT to be more helpful in their daily life, therefore, the “GPT builders” subforum is designed for developers who are building GPTs. Finally, the “ChatGPT” subforum concerns questions or discussions about ChatGPT. Since posts in the remaining subforum (i.e., Community, Announcements, Documentation, and Forum feedback) rarely discuss the challenges faced by LLM developers, we ignore these posts in our empirical study. Specifically, posts in the “Announcements” subforum typically contain official product updates, posts in the “Documentation” subforum offer tutorials and instructional content, posts in the “Community” subforum involve project and technology sharing, and posts in the “Forum feedback” subforum focus on improving the developer forum.
@@
Fig. 2 shows a screenshot of a post [^5] https://community.openai.com/t/547685 from the OpenAI developer forum. Typically, a post consists of several key elements: title, category, tags, question description, code snippet, and the accepted answer (These elements have been marked in this figure). These posts also include additional relevant information, such as creation time, reply time, reply count, view count, and the number of participating users. In this step, we gather the following metadata: the post’s title, creation date, the reply time, and the number of replies. We also collect user metadata including usernames and registration times. In summary, by June 2024, we had crawled data from 29,057 posts and 722,389 users.
Figure 2. A question post from the OpenAI developer forum.
@@
Step 2: Determining popularity trend. To analyze the popularity trend of LLM development among developers, we first perform the time series analysis. Specifically, we count the number of new posts and new users over different periods (i.e., every three months) from the forum’s creation in February 2021 until our data collection deadline in June 2024. Detailed result analysis can be found in Section 4.
Step 3: Determining difficulty level. Questions related to LLM development often present significant challenges when attempting to solve them. To better assess the difficulty of these questions, we analyze them from three key perspectives by following the previous studies on Stack Overflow mining (Abdellatif et al., 2020; Rosen and Shihab, 2016; Alshangiti et al., 2019): reply numbers, accepted solution ratio, and the time required to receive the first reply. Specifically, the number of replies helps gauge community engagement, the accepted solution ratio indicates how frequently a question reaches a satisfactory resolution, and the time required to receive the first reply provides insight into how quickly developers engage with the question. By combining these perspectives, we aim to capture a more holistic view of the difficulty level of LLM-related questions. Detailed result analysis can be found in Section 5.
@@
Step 4: Constructing the taxonomy of challenges. From the 29,057 questions, we randomly selected a statistically significant sample to avoid high cost of manually analyzing all the crawled posts. This sample size ensures a 99% confidence level with a $\pm$2.5confidence interval (Aghajani et al., 2019). As a result, 2,364 questions constitute the taxonomy dataset of our study. We show the taxonomy construction methodology as follows.
@@
First, we randomly select 70% of the 2,364 posts to construct the initial taxonomy of challenges by following the previous study (Wen et al., 2021). This process involves two annotators working in a collaborative way. We use the open coding procedure (Seaman, 1999) to analyze the sampled posts, summarize the preliminary taxonomy, and construct a multi-level hierarchical taxonomy of challenges faced by OpenAI developers. Each annotator has two years of experience in OpenAI development and three years of experience in developing large language models. In this procedure, they need to carefully read the questions of the sampled posts to understand the specific challenges faced by OpenAI developers and the corresponding replies. Specifically, the open coding procedure is conducted as follows: the two annotators read all the sampled posts, with each post being read at least twice. During the analysis, they examine all the data contained in each post, including titles, problem descriptions, code snippets, and replies, summarizing the specific challenges discussed. Then they summarize the specific challenges discussed in the posts. Specifically, some posts describe issues encountered during the use of ChatGPT, for example, “Can’t log in to Chat [^6] https://community.openai.com/t/23899, which are not closely related to developers. In addition, there are also some questions related to account problems, tips for users, etc. These posts are labeled as “Unrelated”. Since the OpenAI developer forum is a platform for all OpenAI users to communicate, some posts are official announcements, document introductions, project discussions, and project sharing. For example, “I have an idea about creating a mix of emoji [^7] https://community.openai.com/t/22180, these posts are temporarily labeled as “Unrelated”. If a post is labeled as “Unrelated”, it means that these posts are not considered when constructing the taxonomy because they are not closely related to OpenAI developers. For the remaining posts, the annotators use simple phrases to summarize the developer’s challenges during the first reading. Specifically, some problem descriptions are about developers consulting on parameter settings when calling OpenAI’s APIs, for example, “GPT-4-Vision-Preview fidelity/detail-parameter [^8] https://community.openai.com/t/477563, the annotators find that this problem is asked because developers do not know how to set API parameters. Therefore, for these issues, the annotators can temporarily describe them as “API Usage Issue, Parameter Setting”.
@@
Finally, we refine the taxonomy of challenges and conduct a reliability analysis. Specifically, the two annotators independently label the remaining 30% of the posts based on the generated preliminary taxonomy. They label each post as “Unrelated” or assign it to the corresponding category. If some posts can not be categorized, they are temporarily labeled as “Pending”. We use Cohen’s Kappa (k(Cohen, 1960) to measure the consistency between the two annotators during the independent labeling process. The computed $k$ value is 0.812, indicating that the two annotators almost completely agree (Landis and Koch, 1977). This demonstrates the reliability of our open coding procedure. In addition, for posts labeled as “Pending”, we introduce the third arbitrator who discusses and analyzes the posts with the two annotators to determine which category these posts ultimately belong to. If they can not be categorized into the preliminary taxonomy, new categories are added. In this phase, we incorporate five additional leaf categories into the preliminary taxonomy framework. Finally, our constructed taxonomy covers all posts related to OpenAI development, and all participants agree on the classification. Detailed result analysis can be found in Section 6.
4. RQ1: Popularity Trend Analysis